# C.I.T. Campus, Taramani, Chennai-600113 Tamil Nadu, India

### v1

@@ Abstract
We explored the effect of space-time geometry on the trajectory of light rays in the context of a charged, rotating black hole. We derived an analytical expression for the deflection of light rays in Kerr-Newman space-time geometry, using a material medium approach, on non-equatorial plane. From this deflection angle expression it is evident that the charge and rotation of the black hole can affect the light rays’ paths. Additionally, we calculated the refractive index of light rays by treating space-time as a medium. We demonstrated how the rotation parameter and charge influence the refractive index and hence the deflection angle of light rays. For Kerr-Newman geometry, the deflection angle decreases with increasing charge when the rotation parameter is held constant. Conversely, for a constant charge, the deflection angle increases with the rotation parameter for prograde and decreases for retrograde trajectories. Applying both factors results in the deflection angle being lower than that of the Schwarzschild geometry. Non-equatorial study of the deflection angle reveals that it is maximum in the equatorial plane than in the pole. The frame-dragging effects in the Kerr-Newman field were taken into account to calculate the velocity of light rays, leading to the determination of the refractive index in this field geometry. This study concludes that depending on the values of the rotation parameter and charge parameter both prograde and retrograde trajectories coincide, resulting in the conclusion that at some point the frame dragging effect is the same for prograde and retrograde motion. Also, the frame dragging increases towards poles for retrograde trajectories while decreasing for prograde trajectories, and these nontrivial nature results because of the interplay between charge and rotation.
@@ 1 Introduction
The study of light deflection in strong and weak gravitational fields began many decades ago. The confirmation of light deflection in 1919 dyson1920ix was the starting point for developing gravitational lensing theory. The deflection of the light ray can be calculated using different methods. The precise deflection of a light beam, when approaching a gravitational mass, is generally obtained by using the null geodesic through which the beam travels. An alternative method to calculate the deflection of a light beam near a gravitational mass is to assume that the light beam is traveling through a material medium, the refractive index of which is determined by the gravitational field. This method estimates the impact of gravitation on the light beam. This is known as the Material Medium Approach.
@@
Several works have been done using null geodesic methods or the Hamilton Jacobi method for the Schwar-zschild black hole (SBH) misner1973gravitation ; weinberg1972principles ; schneider1992gravitational ; virbhadra2000schwarzschild ; keeton2005formalism ; iyer2007light , the Reissner Nordström black hole (RNBH) claudel2001geometry ; eiroa2002reissner ; bozza2002gravitational ; bhadra2003gravitational ; amore2006analytical ; chakraborty2014solar , the Kerr black hole (KBH) bray1986kerr ; kopeikin2002gravitomagnetic ; bozza2003quasiequatorial ; bozza2005analytic ; iyer2009light ; aazami2011alensing ; aazami2011blensing ; chakraborty2017trajectory , and the Kerr-Newman black hole (KNBH) dadhich1977equatorial ; de2000apparent ; hasse2006morse ; kraniotis2014gravitational ; chakraborty2015light ; hsiao2020equatorial ; tinguely2020optical ; hsieh2021gravitational ; hsieh2021strong ; kraniotis2022curvature ; chen2023radii . Furthermore, the material medium approach is also used to calculate the effect of a rotating body on the polarization of light tamm1924electrodynamics ; balazs1958effect , to study light scattering by gravitational field plebanski1960electromagnetic , to obtain the deflection of an electromagnetic wave by gravitational field under different gravitating masses landau2013classical ; de1971gravitational ; mashhoon1973scattering ; mashhoon1975influence ; fischbach1980second . A group of researchers used the effective refractive index to calculate the gravitational time delay and trajectories of light rays in Schwarzschild geometry and showed that the opto- mechanical analogy of general relativity reproduces the equation of GR and matches with the classical equations evans1986f ; evans1996optical . The same work was continued also for the stationary gravitating mass to obtain the trajectory of the photon and obtained the Gravitational Faraday Rotation ishihara1988gravitational ; sereno2003gravitational ; sereno2004weak ; ye2008gravitational . The exact frame-dragging frequency inside the rotating neutron star was obtained which depends on both the radial distance($r$) and the colatitude($\theta$) chakraborty2014dragging . It is also observed that the normalized frame-dragging value is higher at the pole than at the equator for a particular pulsar.
@@
The conditions for equatorial circular geodesics in Kerr-Newman geometry were obtained in 1976 dadhich1977equatorial . In 2000, it was observed that there do not exist any circular photon orbits of radius $r=r_{\pm}$, $r=r_{g}/2$, or $r=0$ in a Kerr-Newman space time and after studying apparent shadows of various KNBHs in the equatorial plane, it is recognized that the spherical symmetry of the shape is destroyed when rotation is present de2000apparent . Hasse and Pelrick hasse2006morse worked on the lensing of KNBH using Morse theory and showed that such a body formed an infinite number of images. Kraniotis kraniotis2014gravitational ; kraniotis2022curvature derived the analytic solutions of the lens equations to calculate the light deflection angle for KNBH and KNBH (anti) de Sitter. Chakraborty and Sen chakraborty2015light obtained the light deflection angle for KNBH in the equatorial plane in terms of mass, spin and charge. In 2020, the deflection angle due to the KNBH in terms of elliptic integrals was obtained hsiao2020equatorial . In 2021, Hsieh et al. hsieh2021strong ; hsieh2021gravitational studied strong gravitational lensing on the equatorial plane and gravitational time delay by KBH and KNBH. In 2023, the equatorial and non-equatorial spherical photon orbits around KNBHs were also studied chen2023radii .
@@
Bozza bozza2003quasiequatorial studied the quasi-equatorial lensing in the strong field limit for spinning BHs. Aazami et al. aazami2011alensing ; aazami2011blensing calculated the two components of the light bending angle, along the direction of the equatorial plane and perpendicular to the equatorial plane of a KBH in the quasi-equatorial regime. Chakraborty and Sen chakraborty2017trajectory investigated the light ray to be travelling only slightly above the equatorial plane for KBH and obtained the off-equatorial light deflection angle for a slowly rotating body.
@@ 2 Kerr-Newman Black Hole
The Kerr-Newman metric provides the precise solution to Einstein’s Field Equation of GR for describing the gravitational field of a rotating, charged BH with stationary and axial symmetry. In the Boyer Lindquist form, in the ($ct$,$r$,$\theta$,$\phi$) coordinate system, the line element of the Kerr-Newman metric is newman1965metric ; newman1965note ; hsiao2020equatorial
$ds^{2}=$ $\left(\frac{\Delta-\alpha^{2}\sin^{2}\theta}{\Sigma^{2}}\right)c^{2}dt^{2}-\frac{\Sigma^{2}}{\Delta}dr^{2}-\Sigma^{2}d\theta^{2}$
@@
$ds^{2}=$ $\left(1-\frac{r_{g}r-Q^{2}}{\Sigma^{2}}\right)c^{2}dt^{2}-\frac{\Sigma^{2}}{\Delta}dr^{2}-\Sigma^{2}d\theta^{2}$
@@
called as the stationary limit surface or the ergosurface. The $r^{E}_{+}(\theta)$ is the outermost surface and the surface $r^{E}_{-}(\theta)$ lies inside the event horizon. The tangent to the static world line is time like for $r>r^{E}_{+}(\theta)$, while for $r<r^{E}_{+}(\theta)$, the world line is space like. The region between the ergosurface and event horizon $r_{+}<r<r^{E}_{+}(\theta)$ is referred as the ergosphere.
@@ 3.1 Polar $(\theta=0)$
The equation admits three possible roots: two positive and one negative. The negative root is unphysical, while among the two positive roots, one lies inside the BH horizon and the other lies outside. For the photon sphere to be physically relevant, it must lie outside the event horizon. Therefore, the physically meaningful photon sphere radius corresponds to the positive root located outside the horizon and is given by {strip}
$r_{\text{ph}}=M+2\sqrt{M^{2}-\frac{1}{3}\left(\alpha^{2}+2Q^{2}\right)}\cos\left[\frac{1}{3}\cos^{-1}\left(\frac{3\sqrt{3}M\left(M^{2}-\alpha^{2}-Q^{2}\right)}{\left(3M^{2}-\alpha^{2}-2Q^{2}\right)^{3/2}}\right)\right].$ (8)
@@ 4 Refractive Index due to Kerr-Newman Black Hole
$ds^{2}=$ $\left(1-\frac{r_{g}r-Q^{2}}{\Sigma^{2}}+\frac{2\alpha(r_{g}r-Q^{2})\sin^{2}\theta}{\Sigma^{2}}\frac{d\phi}{cdt}\right)c^{2}dt^{2}$
@@
$ds^{2}=$ $\left(1-\frac{r_{g}}{r}+\frac{Q^{2}}{r^{2}}+2\alpha\left(\frac{r_{g}}{r}-\frac{Q^{2}}{r^{2}}\right)\sin^{2}\theta\frac{d\phi}{cdt}\right)c^{2}dt^{2}$
@@
$g_{00}=$ $\frac{(1-\frac{r_{g}^{2}}{16\rho^{2}}+\frac{Q^{2}}{4\rho^{2}})^{2}}{\left[(1+\frac{r_{g}}{4\rho})^{2}-(\frac{Q}{2\rho})^{2}\right]^{2}}$
@@
Thus, the isotropic form of KNBH obtained as {strip}
$ds^{2}=$ $\frac{\left(1-\frac{r_{g}^{2}}{16\rho^{2}}+\frac{Q^{2}}{4\rho^{2}}\right)^{2}+2\alpha\left\{\frac{r_{g}}{\rho}\left(\left(1+\frac{r_{g}}{4\rho}\right)^{2}-\left(\frac{Q}{2\rho}\right)^{2}\right)-\frac{Q^{2}}{\rho^{2}}\right\}\sin^{2}\theta\frac{d\phi}{cdt}}{\left[\left(1+\frac{r_{g}}{4\rho}\right)^{2}-\left(\frac{Q}{2\rho}\right)^{2}\right]^{2}}c^{2}dt^{2}-\left\{\left(1+\frac{r_{g}}{4\rho}\right)^{2}-\left(\frac{Q}{2\rho}\right)^{2}\right\}^{2}$
×$\times\left\{d\rho^{2}+\rho^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right\}.$ (20)
The refractive index of the effective medium can be obtained from the isotropic form of the solution (20) by using the material medium approach. Setting ds = 0, the velocity of light ($v(\rho,\theta,\alpha,Q)$) can be identified from the expression of Eq.(20) as
$v(\rho,\theta,\alpha,Q)=\frac{\sqrt{(1-\frac{r_{g}^{2}}{16\rho^{2}}+\frac{Q^{2}}{4\rho^{2}})^{2}+2\alpha\left\{\frac{r_{g}}{\rho}\left((1+\frac{r_{g}}{4\rho})^{2}-(\frac{Q}{2\rho})^{2}\right)-\frac{Q^{2}}{\rho^{2}}\right\}\sin^{2}\theta\frac{d\phi}{cdt}}}{\left[(1+\frac{r_{g}}{4\rho})^{2}-(\frac{Q}{2\rho})^{2}\right]^{2}}.$ (21)
However, this expression of the velocity of light is in the unit of length $\rho$ per unit of time. So, in order to express the velocity in terms of $r$ per unit time, we use Eqs. (15) and (21) and write $v(r,\theta,\alpha,Q)$ as
$\begin{split}v(r,\theta,\alpha,Q)=&v(\rho,\theta,\alpha,Q)\frac{dr}{d\rho}\\ =&\frac{\sqrt{(1-\frac{r_{g}^{2}}{16\rho^{2}}+\frac{Q^{2}}{4\rho^{2}})^{2}+2\alpha\left\{\frac{r_{g}}{\rho}\left((1+\frac{r_{g}}{4\rho})^{2}-(\frac{Q}{2\rho})^{2}\right)-\frac{Q^{2}}{\rho^{2}}\right\}\sin^{2}\theta\frac{d\phi}{cdt}}}{\left[(1+\frac{r_{g}}{4\rho})^{2}-(\frac{Q}{2\rho})^{2}\right]^{2}}\left\{1-\frac{r_{g}^{2}}{16\rho^{2}}+\frac{Q^{2}}{4\rho^{2}}\right\}c\\ =&\left\{\frac{(4\rho)^{2}-(r_{g})^{2}+(2Q)^{2}}{(4\rho+r_{g})^{2}-(2Q)^{2}}\right\}^{2}\sqrt{1+32\alpha\rho r_{g}\frac{(4\rho+r_{g})^{2}-(2Q)^{2}-(4Q)^{2}}{\left\{(4\rho)^{2}-(r_{g})^{2}+(2Q)^{2}\right\}^{2}}\sin^{2}\theta\frac{d\phi}{cdt}}c.\end{split}$ (22)
With value of $4\rho$ from Eq.(13), the velocity $v(r,\theta,\alpha,Q)$ becomes,
$v(r,\theta,\alpha,Q)=$$\left\{1-\frac{r_{g}}{r}+\frac{Q^{2}}{r^{2}}\right\}\sqrt{1+\frac{2(r_{g}r-Q^{2})}{r^{2}-r_{g}r+Q^{2}}\alpha\sin^{2}\theta\frac{d\phi}{cdt}}\,c.$ (23)
Thus the refractive index of the effective medium on non-equatorial plane is
$n(r,\theta,\alpha,Q)=$$\left\{1-\frac{r_{g}}{r}+\frac{Q^{2}}{r^{2}}\right\}^{-1}\left[1+\frac{2(r_{g}r-Q^{2})}{r^{2}-r_{g}r+Q^{2}}\alpha\sin^{2}\theta\frac{d\phi}{cdt}\right]^{-1/2}.$ (24)
Now, by replacing $\frac{r}{r_{g}}$ by $x$ and $\frac{Q}{r_{g}}$ by $q$, we can write the above equation of the refractive index $n(r,\theta,\alpha,Q)$ as follows,
$\begin{split}n(r,\theta,\alpha,Q)=&\left\{1-\frac{1}{x}+\frac{q^{2}}{x^{2}}\right\}^{-1}\left[1+\frac{2(x-q^{2})}{x^{2}-x+q^{2}}\alpha\sin^{2}\theta\frac{d\phi}{cdt}\right]^{-1/2}\\ =&(1-\frac{1}{x})^{-1}\left\{1+\frac{q^{2}}{x(x-1)}\right\}^{-1}\left[1+\frac{2(x-q^{2})}{x^{2}-x+q^{2}}\alpha\sin^{2}\theta\frac{d\phi}{cdt}\right]^{-1/2}\\ =&\left(\frac{x}{x-1}\right)\left\{1+\frac{q^{2}}{x(x-1)}\right\}^{-1}\left[1+\frac{2(x-q^{2})}{x^{2}-x+q^{2}}\alpha\sin^{2}\theta\frac{d\phi}{cdt}\right]^{-1/2}.\end{split}$ (25)
Setting $\alpha=0$ and $Q=0$, we find that the central gravitational mass is static and uncharged, and in such case the above expression of the refractive index goes over to that of Schwarzschild mass which is the same as the refractive index calculated by Sensen2010more for a static non-rotating mass (Schwarzschild geometry). Thus, the second term is in parentheses and the third term is inside the square bracket of the RHS of Eq. (25) is due to the rotation of the charged gravitating body. The second term consists of the charge parameter $q$ and the third term consists of the rotation parameter $\alpha$ and the frame-dragging parameter $\frac{d\phi}{cdt}$ due to the rotating body. For a static uncharged body, these terms just vanish. However, for a charged and very slow rotating body, the RHS of Eq. (25) can be expanded as an infinite converging series, since the second and third terms are significantly less than 1. It may be noted here that Sen’ssen2010more work is restricted to static uncharged geometry. But Eq. (25) can include the case of a charged rotating body (stationary geometry). Therefore, the case discussed by Sensen2010more , appears as a special case when $\alpha$ and $Q$ are set to zero.

### v2

@@ Abstract
We explored the effect of space-time geometry on the trajectory of light rays in the context of a charged, rotating black hole. We derived an analytical expression for the deflection of light rays in Kerr-Newman space-time geometry, using a material medium approach, on non-equatorial plane. From this deflection angle expression it is evident that the charge and rotation of the black hole can affect the light rays’ paths. For Kerr-Newman geometry, the deflection angle decreases with increasing charge when the rotation parameter is held constant. Conversely, for a constant charge, the deflection angle increases with the rotation parameter for prograde and decreases for retrograde trajectories. Applying both factors results in the deflection angle being lower than that of the Schwarzschild geometry. Non-equatorial study of the deflection angle reveals that it is maximum in the equatorial plane than in the pole. The frame-dragging effects in the Kerr-Newman field were taken into account to calculate the velocity of light rays, leading to the determination of the refractive index in this field geometry. This study concludes that, depending on the values of the rotation parameter and charge parameter, both prograde and retrograde trajectories coincide, result in that at some point the frame dragging effect is the same for prograde and retrograde motion. Also, the frame dragging effect increases towards poles than the equatorial plane, and this nontrivial nature results because of the interplay between charge, rotation parameter and rotation frequency of the black hole.
@@ 1 Introduction
The study of light deflection in strong and weak gravitational fields began many decades ago. The confirmation of light deflection in 1919 dyson1920ix was the starting point for developing gravitational lensing theory. The deflection of the light ray can be calculated using different methods. The precise deflection of a light beam, when approaching a gravitational mass, is generally obtained by using the null geodesic through which the beam travels. An alternative method to calculate the deflection of a light beam near a gravitational mass is to assume that the light beam is traveling through a material medium, the refractive index of which is determined by the gravitational field. This method estimates the impact of gravitation on the light beam. This is known as the Material Medium Approach. In this approach, as the light ray propagates through a gravitational field, the effect of gravitation can be equivalently expressed as if the ray of light is passing through a material medium with some specific refractive index. For a centrally symmetric gravitational field, one can obtain a graded refractive index as a function of the radial coordinate $r$ sen2010more . This method can also be extended to an axially symmetric field roy2015trajectory , and in principle to other geometries roy2017deflection ; roy2025deflection . The concept of equivalent material medium was first pointed out by Tamm tamm1924electrodynamics in 1924. Later, this concept was used by Balaz balazs1958effect as early as 1958, and by Plebanski plebanski1960electromagnetic in 1958. Atkinson atkinson1965light investigated the allowed trajectories of light rays near a massive star and showed that the light ray is traveling both radially and tangentially to the field. And a general procedure for utilizing this concept was given by Felice de1971gravitational .
From a scientific standpoint, the material medium approach offers a powerful and intuitive framework for modeling light propagation in curved spacetime, particularly in scenarios involving situations where a number of gravitational bodies are present, such as in binaries or multiple stellar systems, stellar clusters etc. It allows for the incorporation of medium-induced effects on photon trajectories and facilitates comparison with observational signatures, such as angular deflection and time delays. However, it is important to note that this approach does not claim observational superiority over the conventional null geodesic formalism. Both methods are theoretically consistent and independently capable of calculating light deflection. Moreover, the two approaches do not contradict one another; rather than being competing frameworks, they serve as complementary tools, with the choice between them guided by the specific requirements and context of the problem under investigation.
112233445566111.51.5222.52.5rrn(r)n(r)
@@
Several works have been done using null geodesic methods or the Hamilton Jacobi method for the Schwar-zschild black hole (SBH) misner1973gravitation ; weinberg1972principles ; schneider1992gravitational ; virbhadra2000schwarzschild ; keeton2005formalism ; iyer2007light , the Reissner Nordström black hole (RNBH) claudel2001geometry ; eiroa2002reissner ; bozza2002gravitational ; bhadra2003gravitational ; amore2006analytical ; chakraborty2014solar , the Kerr black hole (KBH) bray1986kerr ; kopeikin2002gravitomagnetic ; bozza2003quasiequatorial ; bozza2005analytic ; iyer2009light ; aazami2011alensing ; aazami2011blensing ; chakraborty2017trajectory , and the KNBH dadhich1977equatorial ; de2000apparent ; hasse2006morse ; kraniotis2014gravitational ; chakraborty2015light ; hsiao2020equatorial ; tinguely2020optical ; hsieh2021gravitational ; hsieh2021strong ; kraniotis2022curvature ; Kala:2022uog ; chen2023radii . Furthermore, the material medium approach is also used to calculate the effect of a rotating body on the polarization of light tamm1924electrodynamics ; balazs1958effect , to study light scattering by gravitational field plebanski1960electromagnetic , to obtain the deflection of an electromagnetic wave by gravitational field under different gravitating masses landau2013classical ; de1971gravitational ; mashhoon1973scattering ; mashhoon1975influence ; fischbach1980second . A group of researchers used the effective refractive index to calculate the gravitational time delay and trajectories of light rays in Schwarzschild geometry and showed that the opto- mechanical analogy of general relativity reproduces the equation of GR and matches with the classical equations evans1986f ; evans1996optical . The same work was continued also for the stationary gravitating mass to obtain the trajectory of the photon and obtained the Gravitational Faraday Rotation ishihara1988gravitational ; sereno2003gravitational ; sereno2004weak ; ye2008gravitational . The exact frame-dragging frequency inside the rotating neutron star was obtained which depends on both the radial distance($r$) and the colatitude($\theta$) chakraborty2014dragging . It is also observed that the normalized frame-dragging value is higher at the pole than at the equator for a particular pulsar.
@@
The conditions for equatorial circular geodesics in Kerr-Newman geometry were obtained in 1976 dadhich1977equatorial . In 2000, it was observed that there do not exist any circular photon orbits of radius $r_{\pm}, r=r$$r_{g}/2$, or $r=0$ in a Kerr-Newman space time and after studying apparent shadows of various KNBHs in the equatorial plane, it is recognized that the spherical symmetry of the shape is destroyed when rotation is present de2000apparent . Hasse and Pelrick hasse2006morse worked on the lensing of KNBH using Morse theory and showed that such a body formed an infinite number of images. Kraniotis kraniotis2014gravitational ; kraniotis2022curvature derived the analytic solutions of the lens equations to calculate the light deflection angle for KNBH and KNBH (anti) de Sitter. Chakraborty and Sen chakraborty2015light obtained the light deflection angle for KNBH in the equatorial plane in terms of mass, spin and charge. In 2020, the deflection angle due to the KNBH in terms of elliptic integrals was obtained hsiao2020equatorial . In 2021, Hsieh et al. hsieh2021strong ; hsieh2021gravitational studied strong gravitational lensing on the equatorial plane and gravitational time delay by KBH and KNBH. In 2023, the equatorial and non-equatorial spherical photon orbits around KNBHs were also studied chen2023radii .
@@
Bozza bozza2003quasiequatorial studied the quasi-equatorial lensing in the strong field limit for spinning BHs. Aazami et al. aazami2011alensing ; aazami2011blensing calculated the two components of the light bending angle, along the direction of the equatorial plane and perpendicular to the equatorial plane of a KBH in the quasi-equatorial regime. Chakraborty and Sen chakraborty2017trajectory investigated the light ray to be travelling only slightly above the equatorial plane for KBH and obtained the off-equatorial light deflection angle for a slowly rotating body. Wang et al. Wang:2022ouq derived geodesic trajectories in terms of elliptic integrals and Jacobian elliptic functions for both null and timelike geodesics, expressed as real functions of the Mino time around KNBH spacetimes. Chen et al. Chen:2024oyv studied the extremal KNBH that permits a stable photon orbit exactly at the horizon due to the presence of the degenerate root in the radial potential.
@@ 2 Kerr-Newman Black Hole
[…]
$ds^{2}=\$$\left(\frac{\Delta-\alpha^{2}\sin^{2}\theta}{\Sigma^{2}}\right)c^{2}dt^{2}-\frac{\Sigma^{2}}{\Delta}dr^{2}-\Sigma^{2}d\theta^{2}$
@@
$ds^{2}=\$$\left(1-\frac{r_{g}r-Q^{2}}{\Sigma^{2}}\right)c^{2}dt^{2}-\frac{\Sigma^{2}}{\Delta}dr^{2}-\Sigma^{2}d\theta^{2}$
@@
called the stationary limit surface or the ergosurface. The $r^{E}_{+}(\theta)$ is the outermost surface and the surface $r^{E}_{-}(\theta)$ lies inside the event horizon. The tangent to the static world line is time like for $r>r^{E}_{+}(\theta)$, while for $r<r^{E}_{+}(\theta)$, the world line is space like. The region between the ergosurface and the event horizon $r_{+}<r<r^{E}_{+}(\theta)$ is referred as the ergosphere.
@@ 3.1 Polar $(\theta=0)$
The equation admits three possible roots: two positive and one negative. The negative root is unphysical, while among the two positive roots, one lies inside the BH horizon and the other lies outside. For the photon sphere to be physically relevant, it must lie outside the event horizon. Therefore, the physically meaningful photon sphere radius corresponds to the positive root located outside the horizon and is given by
$\begin{split}r_{\text{ph}}=\ &M+2\sqrt{M^{2}-\frac{1}{3}\left(\alpha^{2}+2Q^{2}\right)}\\ &\times\cos\left[\frac{1}{3}\cos^{-1}\left(\frac{3\sqrt{3}M\left(M^{2}-\alpha^{2}-Q^{2}\right)}{\left(3M^{2}-\alpha^{2}-2Q^{2}\right)^{3/2}}\right)\right].\end{split}$ (8)
@@ 4 Refractive Index due to Kerr-Newman Black Hole
$ds^{2}=\$$\left(1-\frac{r_{g}r-Q^{2}}{\Sigma^{2}}+\frac{2\alpha(r_{g}r-Q^{2})\sin^{2}\theta}{\Sigma^{2}}\frac{d\phi}{cdt}\right)c^{2}dt^{2}$
@@
$ds^{2}=\$$\left(1-\frac{r_{g}}{r}+\frac{Q^{2}}{r^{2}}+2\alpha\left(\frac{r_{g}}{r}-\frac{Q^{2}}{r^{2}}\right)\sin^{2}\theta\frac{d\phi}{cdt}\right)c^{2}dt^{2}$
@@
$g_{00}=\$$\frac{(1-\frac{r_{g}^{2}}{16\rho^{2}}+\frac{Q^{2}}{4\rho^{2}})^{2}}{\left[(1+\frac{r_{g}}{4\rho})^{2}-(\frac{Q}{2\rho})^{2}\right]^{2}}$
@@
Thus, the isotropic form of KNBH obtained as
Setting $\alpha=0$ and $Q=0$, we find that the central gravitational mass is static and uncharged, and in such case the above expression of the refractive index goes over to that of Schwarzschild mass which is the same as the refractive index calculated by Sensen2010more for a static non-rotating mass (Schwarzschild geometry). Thus, the second term is in parentheses and the third term is inside the square bracket of the RHS of Eq. (LABEL:Eq_ref_x) is due to the rotation of the charged gravitating body. The second term consists of the charge parameter $q$ and the third term consists of the rotation parameter $\alpha$ and the frame-dragging parameter $\frac{d\phi}{cdt}$ due to the rotating body. For a static uncharged body, these terms just vanish. However, for a charged and very slow rotating body, the RHS of Eq. (LABEL:Eq_ref_x) can be expanded as an infinite converging series, since the second and third terms are significantly less than 1. It may be noted here that Sen’ssen2010more work is restricted to static uncharged geometry. But Eq. (LABEL:Eq_ref_x) can include the case of a charged rotating body (stationary geometry). Therefore, the case discussed by Sensen2010more , appears as a special case when $\alpha$ and $Q$ are set to zero.